# Quantifying interdisciplinary synergy in higher STEM education

### v2

@@ Abstract
We propose a framework to quantify and utilize interdisciplinarity in science and engineering curricula in the case of university-level higher education. We analyze interdisciplinary relations by standardizing large-scale official educational data in Korea using a cutting-edge large language model and constructing knowledge maps for disciplines of scientific education. We design and evaluate single-field and integrated dual-field curricula by adapting pedagogical theory and utilizing information theory-based metrics. We develop standard curricula for individual disciplines and integrated curricula combining two fields, with their interdisciplinarity quantified by the curriculum synergy score. The results indicate higher interdisciplinarity for combinations within or across closely related fields, especially in engineering fields. Based on the analysis, engineering fields constitute the core structure of our design for curriculum interdisciplinarity, while basic natural science fields are located at peripheral stems to provide fundamental concepts.
@@ 1 Introduction
The increasing complexity of challenges across the social, academic, and industrial domains has focused on interdisciplinary approaches to solving advanced problems [1]. In response to this social demand, higher education institutions, e.g., universities, are progressively introducing and expanding their multidisciplinary programs [2, 3, 4, 5]. Consequently, researchers are actively working on the interdisciplinary theory and its curriculum models nowadays [2, 6]. Unfortunately, however, most of such research relies on qualitative assessments based on theoretical arguments or small-scale data, which makes its versatile application questionable.
@@
The paper is organized as follows. In Sec. 2, we introduce our data and the process of course standardization with the large language model (LLM). Based on the processed data, we first analyze the structural properties of the department similarity in terms of curricula in Sec. 3. As the main topic of our work, we present the interdisciplinary curriculum model and the quantitative measure of synergy between disciplines in Sec. 4. We summarize our work and discuss its implications in Sec. 5.
@@ 2 Data processing of course standardization
As introduced in Sec. 1, we focus on the departments running four-year undergraduate programs within STEM fields classified under Engineering and Natural Sciences in the Standard Classification (the highest-level classification). Specifically, within Natural Sciences, we extracted data from departments categorized as “Basic Natural Sciences,” which include “Chemistry, Life Science, and Environment” and “Mathematics, Physics, Astronomy, and Earth Science” in the Standard Classification (the intermediate-level classification). As a result of our preprocessing and comprehensive data integration, we constructed a standardized dataset comprising curricula from 2 84128412\,8412 841 STEM departments [^1] The original data included 2 85128512\,8512 851 unique STEM undergraduate departments. However, due to military security restrictions and data structure errors, only 2 84128412\,8412 841 departments were ultimately analyzed. across $161$ higher education institutions, including a total of 126 437126437126\,437126 437 raw course entries. We use departments as the primary unit of analysis because our data, sourced from the Ministry of Education in Korea, is organized at this level, providing a consistent framework with around $80$ credits (for major courses) over four years per department. In Korea, departments serve as the de facto units of disciplines due to the structured nature of higher education, aligning with current educational practices and government initiatives that encourage interdisciplinary integration. While finer distinctions at the program level might be desirable, current data limitations necessitate this approach. Future research could explore more granular analyses as detailed data becomes available, such as through “microdegree” programs aimed at fostering small units of interdisciplinary education.
@@
The courses are obviously building blocks of the coursework in each department. However, we cannot simply use the raw course data as it is, because there are multiple nomenclatural variants of essentially the same course, e.g., “quantum mechanics” and “quantum physics” in physics-related departments. Therefore, it is crucial to standardize different variants of course names in the raw data to accurately identify course units that are as meaningfully independent as possible shared by different departments and institutions. Without the process, only the courses with exactly the same name would be identified as the same entity. Such a task in this scale would be next to impossible manually, but cutting-edge LLMs enable us to do this task systematically. In particular, we use Gemini-1.5 Pro [7], which is commonly referred to as one of the most accurate language models as of July 2024, when we performed this part of data processing. By employing higher-order data processing techniques combined with the LLM, which will be discussed in Sec. 2.2, we systematically process these raw courses into 24 4762447624\,47624 476 standardized courses. Our approach enables us to effectively process large-scale datasets while accurately capturing the semantic complexities inherent in natural language and disciplinary knowledge.
Figure 1: A schematic illustration of education data processing with LLM. (a) The tokenizing process of course data by LLM (Sec. 2.1). Each course is broken into distinct tokens, marked by different colors. (b) For each token, the courses including the token are listed below. Among them, the courses identified as the same by the LLM are connected in the form of hyperedges (enclosed by the same connected shaded area). (c) Each connected component constructed by collecting the hyperedges in the process described in the panel (b) is identified as each standardized course.
@@ 2.1 Tokenizing from the raw course data with LLM
Even with the LLM, extracting the standardized courses directly from the exhaustive list of 126 437126437126\,437126 437 raw courses is not plausible due to the combinatorial explosion. Our strategy to circumvent this problem to proceed is the following. We first utilize tokenization to segment each of the raw course names into meaningful components that serve as the basic indicators for further processing. Each course $C_{i}$ is decomposed into a set of tokens $\{t_{1},t_{2},\dots,t_{m}\}$ as illustrated in Fig. 1(a). To ensure accurate processing of the natural language in the raw course names (Korean), we use prompts tailored to its linguistic properties with specific examples (prompt 1 in Appendix A). Through this tokenization process, we obtain 8 84188418\,8418 841 unique tokens from the raw course data. By focusing on the shared tokens among different courses, this preprocesses the data to highlight potential semantic similarities for further analysis. To process raw courses effectively, for each token, we collect the list of courses including it as illustrated in Fig. 1(b).
@@ 2.2 Constructing hyperedges among raw courses with LLM
Now we have the set of raw courses for each token $t$ as in Fig. 1(b), and our next step is to collect the effectively same raw courses as intermediate-level units. Mathematically, this step involves identifying higher-order relationships (i.e., hyperedges) [8] among raw courses—beyond simple pairwise associations—so as to capture more nuanced semantic groupings that reflect disciplinary structures [16, 17, 18]. Similar to the tokenization process described in Sec. 2.1, we again use a prompt that effectively clusters identical courses based on shared topics or subfields (prompt 2 in Appendix A), depicted in Fig. 1(b).
@@
Through this LLM-based higher-order clustering process, we generate 91 9649196491\,96491 964 hyperedges from all tokens, which encapsulate the intermediate-level interrelationships present in the raw course data.
@@ 2.3 Standardizing courses
Our final step to standardize courses is based on the list of hyperedges in Sec. 2.2. As illustrated in Fig. 1(c), we merge all of the courses inside each of the connected components. Intuitively, this definition captures not only directly connected courses within a hyperedge but also courses that are indirectly linked through the chain of associations. For example, if course $C_{1}$ is associated with hyperedge $H_{k}$ and course $C_{2}$ is linked to both $H_{k}$ and $H_{l}$, then $C_{1}$ and $C_{2}$ belong to the same connected component. The set of connected components constitutes our final set of standardized courses $\mathcal{S}$. In this way, each connected component groups the raw courses that are semantically or structurally similar based on shared tokens or associations identified via the hyperedge construction process.
Through this standardization process, we reduce 126 347126347126\,347126 347 raw courses to 24 4762447624\,47624 476 standardized courses from STEM departments of Korean universities. This reduction merges redundant or overlapping courses into unified representations while preserving their semantic integrity and higher-order relationships, for more interpretable and robust analysis. By our processing, the standardized courses provide a deeper understanding of the underlying structures and interconnections within curricula [20]. For a concrete example, we provide the step-by-step process in the case of statistical physics in Appendix B.
@@ 3 Department similarity based on standardized courses
Figure 3: Alluvial diagram showing the membership reorganization of community structures at four representative resolution values: $\gamma=0$, $0.56$, $1.12$, and $1.77$, where the community structures show global consistency (see Appendix C for details), where the vertical positions of the nodes are determined to minimize crossing. At $\gamma=0.56$, each node and link represents the community’s aggregated departmental leanings (PSE: red, TSE: green, and MLSE: blue), and at finer resolution values, the weighted sum of colors represents the reorganization of departmental classifications.
To verify the validity and utility of the standardized courses extracted in Sec. 2, we constructed a department similarity network (DSN), where nodes represent departments, and the weighted edges represent their department similarity in terms of their curricula (Fig. 2[21, 22]. The department similarity is quantified using the weighted Jaccard similarity coefficient, defined as:
@@
We explore the structural properties of DSN, in the viewpoint of communities [23, 24] of various scales. We take the popular modularity-maximization algorithm Louvain [25] with the tunable resolution parameter $\gamma$. We choose four representative values of $\gamma=0$, $0.56$, $1.12$, and $1.77$, where the overall community structures show global consistency among the results (Appendix C[26]. We illustrate the scale-dependent reorganization of communities across the four resolution values as an alluvial diagram in Fig. 3. Each block in the diagram corresponds to a community within DSN at a particular $\gamma$ value, with the inter-connection between the blocks in the adjacent $\gamma$ values indicating the flow of department membership.
@@ 4 Interdisciplinary curriculum design and quantifying its synergy
Figure 4: A schematic diagram of our curriculum design process. (a) A recap of the weighted bipartite network between departments and standardized courses, described as a weighted matrix in Fig. 2. The detailed course standardization process is stated in Sec. 2. (b) The standardized-course-mode projection from the bipartite network, from which the curriculum is aggregated by sequential addition of candidate courses based on the information-theoretical criterion described in panel (c).
In this section, based on the standardized courses, we suggest a framework for designing curricula for potential interdisciplinary fields, followed by evaluating the synergy from such field combinations. First, we need to set up a list of standard academic disciplines. As a reminder, we have constructed a comprehensive list of STEM curricula in terms of standardized courses for each department in Sec. 2 and Sec. 3. This process involves constructing a bipartite network with nodes representing 24 4672446724\,46724 467 standardized courses and 2 84128412\,8412 841 STEM departments. The edges in the bipartite network are weighted by the total credit hours associated with each course [Figs. 2 and 4(a)]. To quantify the relationship between the standardized courses, we project the bipartite network into the unipartite network composed of standardized courses, whose edges are weighted by the sum of credits in the departments sharing the course pairs [Fig. 4(b)].
@@ 4.1 Single-discipline curriculum model
From the original source data provided by Ministry of Education in Korea [15], the classification of “Standard Classification information (the lowest-level classification)” is designated as the list of $51$ independent disciplines [^2] Among the $52$ disciplines in the data set [15], we exclude the “not classified elsewhere (N.C.E.)” entry for proper analysis. . For each discipline, we define the standardized credits of the standardized courses in the discipline as the most frequent nonzero credits in the departments belonging to the discipline. For instance, if the credits of a standardized course $\mathcal{C}$ appear as $\{6,6,3,6,4,0,0,0,0,0\}$ in $10$ departments corresponding to a discipline $\mathcal{D}$, the standardized credit of $\mathcal{C}$ in discipline $\mathcal{D}$ is $6$. From now on, “courses” and “credits” always refer to standardized ones, unless stated otherwise.
@@
The addition of the candidate course in the model is accepted if $\langle\mathcal{J}\rangle_{\mathcal{D}}\leq 0$, and rejected if $\langle\mathcal{J}\rangle_{\mathcal{D}}>0$. Finally, we terminate the model aggregation when the sum of the credits reaches at least $80$, which is a typical number of credits for major courses in four-year undergraduate degrees in Korea. The optimized curriculum of discipline $\mathcal{D}$, denoted by $\Omega\left(\mathcal{D}\right)$, can be expressed as the credit-weighted vector for each course
@@
where $\tilde{w}_{\mathcal{D};i}$ is the credit of course $i$ in the optimized curriculum of discipline $\mathcal{D}$. The vector for each discipline is standardized by the optimization process described in Sec. 2, resulting in the uniform credit sum $\approx 80$, to ensure comparability and consistency in the curriculum design process.
@@ 4.2 Interdisciplinary curriculum model
To construct a double-field interdisciplinary curriculum model, we start with a seed of five courses in each field and their assigned credits for the $10$ courses in total. Similar to the single-discipline model in Sec. 4.1, the candidate courses for addition are interactively sorted by the sum of weights [in the projected network in Fig. 4(b)] attached to the current courses in the model.
@@
This iterative process retains only the courses that reduce the average JSD across all of the departments belonging to both disciplines, and the same threshold of $80$ credits applies for the model termination. Similar to the single-discipline case, the optimized interdisciplinary curriculum of disciplines $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, denoted by $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$, can be expressed as the credit-weighted vector for each course
@@ 4.3 Quantifying interdisciplinarity with the curriculum synergy score
In Sec. 4.2, we have proposed an information-theoretical model to evaluate the degree of integration achieved by interdisciplinary curricula. The central aim is to design a curriculum that reflects both semantic cohesion across departments and pedagogical completeness within a combined field of study. To formalize this, we define an optimized curriculum set $\Omega(\mathcal{D})$ for any given set of departments $\mathcal{D}$. The selection process is based on a semantic filtering criterion: each course must contribute to reducing the average divergence in Eq. (5) across departmental curricula. The credit constraint mentioned in Sec. 4.1 ensures that the resulting curriculum maintains the completeness required for a full major according to national education standards. An analogous formulation is applied to interdisciplinary combinations $\mathcal{D}_{A}\otimes\mathcal{D}_{B}$. Based on this formal construction, we introduce the curriculum synergy score (CSS), which evaluates whether the algorithmically constructed interdisciplinary curriculum $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$ achieves a deeper level of integration than a simple combination of the two disciplines, denoted by $\Omega\left(\mathcal{D}_{A}\cup\mathcal{D}_{B}\right)$.
@@
The CSS, therefore, quantifies the synergy between disciplines that cannot solely be captured by simple similarity measures, which highlights the added value of proper interdisciplinary integration. We observe the distribution of each factor separately, and one can check that the vast majority of $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ values are close to 00 [Fig. 5(a)], which is understandable by the fact that most discipline pairs would be quite different to each other. On the other hand, the distribution of $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ shows the most prominent peak around $0.3$ presumably corresponding to the overall scale difference between $\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}}$ and $\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}$, and the second tallest peak around $0.6$ [Fig. 5(b)]. There is no notable correlation between $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ and $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ [Fig. 5(c)], which also justifies the formalism of $\Theta\left(\mathcal{D}_{A},\mathcal{D}_{B}\right)$ as the product of them. In Appendix D, we provide the plots in the double-logarithmic scale in Fig. 9 to help the reader better understand the range and tail behavior of the similarity values.
For notable examples, we provide three examples in supplementary tables: two examples of beneficial synergy (‘physics’ + ‘semiconductor engineering’ in Table S1 and ‘chemistry’ + ‘high polymer engineering’ in Table S2) and an example of no synergy (‘physics’ + ‘chemistry’ in Table S3). It is interesting to note that science disciplines and their related engineering disciplines are much better combinations than two science disciplines, and this tendency will be highlighted again in the analysis presented in the next subsection as well. Other examples include combinations ‘biotechnology + energy engineering,’ which are intuitively synergistic, alongside less conventional yet plausible pairs such as ‘nautical science + civil engineering’ and ‘optical Engineering + semiconductor engineering.’

### v3

@@ Abstract
We propose a framework to quantify and utilize interdisciplinarity in science and engineering curricula at the university-level higher education. We analyze interdisciplinary relations by standardizing large-scale official educational data in Korea using a cutting-edge large language model and constructing knowledge maps for disciplines of scientific education. We design and evaluate single-field and integrated dual-field curricula by adapting pedagogical theory and utilizing information theory-based metrics. We develop standard curricula for individual disciplines and integrated curricula combining two fields, with their interdisciplinarity quantified by the curriculum synergy score. The results indicate higher interdisciplinarity for combinations within or across closely related fields, especially in engineering fields. Based on the analysis, engineering fields constitute the core structure of our design for curriculum interdisciplinarity, while basic natural science fields are located at peripheral stems to provide fundamental concepts.
@@ 1 Introduction
The increasing complexity of challenges across the social, academic, and industrial domains has focused on interdisciplinary approaches to solving advanced problems [1]. In response to this social demand, higher education institutions, e.g., universities, are progressively introducing and expanding their multidisciplinary programs [2, 3, 4, 5]. Consequently, researchers are actively working on the interdisciplinary theory and its curriculum models nowadays [2, 6]. Unfortunately, most such research relies on qualitative assessments based on theoretical arguments or small-scale data, calling into question its generalizability.
@@
The paper is organized as follows. In Sect. 2, we introduce our data and the process of course standardization with the large language model (LLM). Based on the processed data, we first analyze the structural properties of the department similarity in terms of curricula in Sect. 3. As the main topic of our work, we present the interdisciplinary curriculum model and the quantitative measure of synergy between disciplines in Sect. 4. We summarize our work and discuss its implications in Sect. 5.
@@ 2 Data processing of course standardization
As introduced in Sect. 1, we focus on the departments running four-year undergraduate programs within STEM fields classified under Engineering and Natural Sciences in the Standard Classification (the highest-level classification). Specifically, within Natural Sciences, we extracted data from departments categorized as “Basic Natural Sciences,” which include “Chemistry, Life Science, and Environment” and “Mathematics, Physics, Astronomy, and Earth Science” in the Standard Classification (the intermediate-level classification). As a result of our preprocessing and comprehensive data integration, we constructed a standardized dataset comprising curricula from 2 8412\,8412 841 STEM departments [^1] The original data included 2 8512\,8512 851 unique STEM undergraduate departments. However, due to military security restrictions and data structure errors, only 2 8412\,8412 841 departments were ultimately analyzed. across 161161161 higher education institutions, including a total of 126 437126\,437126 437 raw course entries. We use departments as the primary unit of analysis because our data, sourced from the Ministry of Education in Korea, is organized at this level, providing a consistent framework with around 808080 credits (for major courses) over four years per department. In Korea, departments serve as the de facto units of disciplines due to the structured nature of higher education, aligning with current educational practices and government initiatives that encourage interdisciplinary integration. While finer distinctions at the program level might be desirable, current data limitations necessitate this approach. Future research could explore more granular analyses as detailed data becomes available, such as through “microdegree” programs aimed at fostering small units of interdisciplinary education.
@@
The courses are obviously building blocks of the coursework in each department. However, we cannot simply use the raw course data as it is, because there are multiple nomenclatural variants of essentially the same course, e.g., “quantum mechanics” and “quantum physics” in physics-related departments. Therefore, it is crucial to standardize different variants of course names in the raw data to accurately identify course units that are as meaningfully independent as possible shared by different departments and institutions. Without the process, only the courses with exactly the same name would be identified as the same entity. Such a task in this scale would be next to impossible manually, but cutting-edge LLMs enable us to do this task systematically. In particular, we use Gemini-1.5 Pro [7], which is commonly referred to as one of the most accurate language models as of July 2024, when we performed this part of data processing. By employing higher-order data processing techniques combined with the LLM, which will be discussed in Sect. 2.2, we systematically process these raw courses into 24 47624\,47624 476 standardized courses. Our approach enables us to effectively process large-scale datasets while accurately capturing the semantic complexities inherent in natural language and disciplinary knowledge.
Figure 1: A schematic illustration of education data processing with LLM. (a) The tokenizing process of course data by LLM (Sect. 2.1). Each course is broken into distinct tokens, marked by different colors. (b) For each token, the courses including the token are listed below. Among them, the courses identified as the same by the LLM are connected in the form of hyperedges (enclosed by the same connected shaded area). (c) Each connected component constructed by collecting the hyperedges in the process described in the panel (b) is identified as each standardized course.
@@ 2.1 Tokenizing from the raw course data with LLM
Even with the LLM, extracting the standardized courses directly from the exhaustive list of 126 437126\,437126 437 raw courses is not plausible due to the combinatorial explosion. Our strategy to circumvent this problem to proceed is the following. We first utilize tokenization to segment each of the raw course names into meaningful components that serve as the basic indicators for further processing. Each course $C_{i}$ is decomposed into a set of tokens $\{t_{1},t_{2},\dots,t_{m}\}$ as illustrated in Fig. 1(a). To ensure accurate processing of the natural language in the raw course names (Korean), we use prompts tailored to its linguistic properties with specific examples (prompt 1 in Appendix A). Through this tokenization process, we obtain 8 8418\,8418 841 unique tokens from the raw course data. By focusing on the shared tokens among different courses, this preprocesses the data to highlight potential semantic similarities for further analysis. To process raw courses effectively, for each token, we collect the list of courses including it as illustrated in Fig. 1(b).
@@ 2.2 Constructing hyperedges among raw courses with LLM
Now we have the set of raw courses for each token $t$ as in Fig. 1(b), and our next step is to collect the effectively same raw courses as intermediate-level units. Mathematically, this step involves identifying higher-order relationships (i.e., hyperedges) [8] among raw courses—beyond simple pairwise associations—so as to capture more nuanced semantic groupings that reflect disciplinary structures [16, 17, 18]. Similar to the tokenization process described in Sect. 2.1, we again use a prompt that effectively clusters identical courses based on shared topics or subfields (prompt 2 in Appendix A), depicted in Fig. 1(b).
@@
Through this LLM-based higher-order clustering process, we generate 91 96491\,96491 964 hyperedges from all tokens, which encapsulate the intermediate-level interrelationships present in the raw course data.
@@ 2.3 Standardizing courses
Our final step to standardize courses is based on the list of hyperedges in Sect. 2.2. As illustrated in Fig. 1(c), we merge all of the courses inside each of the connected components. Intuitively, this definition captures not only directly connected courses within a hyperedge but also courses that are indirectly linked through the chain of associations. For example, if course $C_{1}$ is associated with hyperedge $H_{k}$ and course $C_{2}$ is linked to both $H_{k}$ and $H_{l}$, then $C_{1}$ and $C_{2}$ belong to the same connected component. The set of connected components constitutes our final set of standardized courses $\mathcal{S}$. In this way, each connected component groups the raw courses that are semantically or structurally similar based on shared tokens or associations identified via the hyperedge construction process.
Through this standardization process, we reduce 126 347126\,347126 347 raw courses to 24 47624\,47624 476 standardized courses from STEM departments of Korean universities. This reduction merges redundant or overlapping courses into unified representations while preserving their semantic integrity and higher-order relationships, for more interpretable and robust analysis. By our processing, the standardized courses provide a deeper understanding of the underlying structures and interconnections within curricula [20]. For a concrete example, we provide the step-by-step process in the case of statistical physics in Appendix B.
@@ 3 Department similarity based on standardized courses
Figure 3: Alluvial diagram showing the membership reorganization of community structures at four representative resolution values: $\gamma=0$, 0.560.560.56, 1.121.121.12, and 1.771.771.77, where the community structures show global consistency (see Appendix C for details), where the vertical positions of the nodes are determined to minimize crossing. At $\gamma=0.56$, each node and link represents the community’s aggregated departmental leanings (PSE: red, TSE: green, and MLSE: blue), and at finer resolution values, the weighted sum of colors represents the reorganization of departmental classifications.
To verify the validity and utility of the standardized courses extracted in Sect. 2, we constructed a department similarity network (DSN), where nodes represent departments, and the weighted edges represent their department similarity in terms of their curricula (Fig. 2[21, 22]. The department similarity is quantified using the weighted Jaccard similarity coefficient, defined as:
@@
We explore the structural properties of DSN from the perspective of communities [23, 24] of various scales. We take the popular modularity-maximization algorithm Louvain [25] with the tunable resolution parameter $\gamma$. We choose four representative values of $\gamma=0$, 0.560.560.56, 1.121.121.12, and 1.771.771.77, where the overall community structures show global consistency among the results (Appendix C[26]. We illustrate the scale-dependent reorganization of communities across the four resolution values as an alluvial diagram in Fig. 3. Each block in the diagram corresponds to a community within DSN at a particular $\gamma$ value, with the inter-connection between the blocks in the adjacent $\gamma$ values indicating the flow of department membership.
@@ 4 Interdisciplinary curriculum design and quantifying its synergy
Figure 4: A schematic diagram of our curriculum design process. (a) A recap of the weighted bipartite network between departments and standardized courses, described as a weighted matrix in Fig. 2. The detailed course standardization process is stated in Sect. 2. (b) The standardized-course-mode projection from the bipartite network, from which the curriculum is aggregated by sequential addition of candidate courses based on the information-theoretical criterion described in panel (c).
In this section, based on the standardized courses, we suggest a framework for designing curricula for potential interdisciplinary fields, followed by evaluating the synergy from such field combinations. First, we need to set up a list of standard academic disciplines. As a reminder, we have constructed a comprehensive list of STEM curricula in terms of standardized courses for each department in Sect. 2 and Sect. 3. This process involves constructing a bipartite network with nodes representing 24 46724\,46724 467 standardized courses and 2 8412\,8412 841 STEM departments. The edges in the bipartite network are weighted by the total credit hours associated with each course [Figs. 2 and 4(a)]. To quantify the relationship between the standardized courses, we project the bipartite network into the unipartite network composed of standardized courses, whose edges are weighted by the sum of credits in the departments sharing the course pairs [Fig. 4(b)].
@@ 4.1 Single-discipline curriculum model
[…]
@@
The addition of the candidate course in the model is accepted if $\langle\mathcal{J}\rangle_{\mathcal{D}}\leq 0$, and rejected if $\langle\mathcal{J}\rangle_{\mathcal{D}}>0$. Finally, we terminate the model aggregation when the sum of the credits reaches at least 808080, which is a typical number of credits for major courses in four-year undergraduate degrees in Korea. The optimized curriculum of discipline $\mathcal{D}$, denoted by $\Omega\left(\mathcal{D}\right)$, can be expressed as the credit-weighted vector for each course
@@
where $\tilde{w}_{\mathcal{D};i}$ is the credit of course $i$ in the optimized curriculum of discipline $\mathcal{D}$. The vector for each discipline is standardized by the optimization process described in Sect. 2, resulting in the uniform credit sum 80$\approx$ 80≈ 80, to ensure comparability and consistency in the curriculum design process.
@@ 4.2 Interdisciplinary curriculum model
To construct a double-field interdisciplinary curriculum model, we start with a seed of five courses in each field and their assigned credits for the 101010 courses in total. Similar to the single-discipline model in Sect. 4.1, the candidate courses for addition are interactively sorted by the sum of weights [in the projected network in Fig. 4(b)] attached to the current courses in the model.
@@
This iterative process retains only the courses that reduce the average JSD across all of the departments belonging to both disciplines, and the same threshold of 808080 credits applies for the model termination. Similar to the single-discipline case, the optimized interdisciplinary curriculum of disciplines $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, denoted by $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$, can be expressed as the credit-weighted vector for each course
@@ 4.3 Quantifying interdisciplinarity with the curriculum synergy score
In Sect. 4.2, we have proposed an information-theoretical model to evaluate the degree of integration achieved by interdisciplinary curricula. The central aim is to design a curriculum that reflects both semantic cohesion across departments and pedagogical completeness within a combined field of study. To formalize this, we define an optimized curriculum set $\Omega(\mathcal{D})$ for any given set of departments $\mathcal{D}$. The selection process is based on a semantic filtering criterion: each course must contribute to reducing the average divergence in Eq. (5) across departmental curricula. The credit constraint mentioned in Sect. 4.1 ensures that the resulting curriculum maintains the completeness required for a full major according to national education standards. An analogous formulation is applied to interdisciplinary combinations $\mathcal{D}_{A}\otimes\mathcal{D}_{B}$. Based on this formal construction, we introduce the curriculum synergy score (CSS), which evaluates whether the algorithmically constructed interdisciplinary curriculum $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$ achieves a deeper level of integration than a simple combination of the two disciplines, denoted by $\Omega\left(\mathcal{D}_{A}\cup\mathcal{D}_{B}\right)$.
@@
The CSS, therefore, quantifies the synergy between disciplines that cannot solely be captured by simple similarity measures, which highlights the added value of proper interdisciplinary integration. We observe the distribution of each factor separately, and one can check that the vast majority of $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ values are close to 0 [Fig. 5(a)], which is understandable by the fact that most discipline pairs would be quite different to each other. On the other hand, the distribution of $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ shows the most prominent peak around 0.30.30.3 presumably corresponding to the overall scale difference between $\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}}$ and $\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}$, and the second tallest peak around 0.60.60.6 [Fig. 5(b)]. There is no notable correlation between $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ and $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ [Fig. 5(c)], which also justifies the formalism of $\Theta\left(\mathcal{D}_{A},\mathcal{D}_{B}\right)$ as the product of them. In Appendix D, we provide the plots in the double-logarithmic scale in Fig. 9 to help the reader better understand the range and tail behavior of the similarity values.
For notable examples, we provide three examples in supplementary tables: two examples of beneficial synergy (‘physics’ + ‘semiconductor engineering’ in Table S1 and ‘chemistry’ + ‘high polymer engineering’ in Table S2) and an example of no synergy (‘physics’ + ‘chemistry’ in Table S3). Notably, science disciplines and their related engineering disciplines form better combinations than two science disciplines, and this tendency will be highlighted again in the analysis presented in the next subsection. Other examples include combinations ‘biotechnology + energy engineering,’ which are intuitively synergistic, alongside less conventional yet plausible pairs such as ‘nautical science + civil engineering’ and ‘optical Engineering + semiconductor engineering.’